OpenStudy (anonymous):

Determine if the subset f(x)=c , the set of all constant functions is a subspace of all continous functions

OpenStudy (turingtest):

sure seems that way to me c+d=e another constant a*c=ac another constant so I think it is closed under addition and multiplication

OpenStudy (turingtest):

or a*c=d however you want to write it...

OpenStudy (anonymous):

okkie dokkie

OpenStudy (anonymous):

I have one more lol

OpenStudy (turingtest):

wait I totally misunderstood that question :/

OpenStudy (anonymous):

ohh lol

OpenStudy (turingtest):

I thought constant functions lol

OpenStudy (turingtest):

I mean, wait... no that is what they mean, right? if constants are a subset of continuous functions?

OpenStudy (anonymous):

i think so that is what i thought

OpenStudy (turingtest):

ok then I'm right after all sorry, other question?

OpenStudy (anonymous):

umm when f(0)=1

OpenStudy (anonymous):

umm that means that c=1

OpenStudy (anonymous):

no?

OpenStudy (turingtest):

talking about the same problem?

OpenStudy (anonymous):

yes like instead of it being f(x)=c it is f(x)=1

OpenStudy (anonymous):

whoops f(0)=1

OpenStudy (turingtest):

f(x)=c it doesn't matter what x is, c is whatever we choose it to be if you mean a different problem, like asking about if f(0)=1 is a subspace of c, the answer is no

OpenStudy (anonymous):

noo ok let me rephrase the question

OpenStudy (anonymous):

Determine if the subset f(0)=1 is a subspace of all continous functions

OpenStudy (anonymous):

now that i am looking at it i think the answer wld be yes

OpenStudy (turingtest):

that is what I meant by if f(0)=1 is a subspace of c ask yourself this: what would be f(0)+g(0) in this vector space ?

OpenStudy (anonymous):

c+d which is a constant which is a straight line. so it wld still be continous

OpenStudy (turingtest):

but\[f(0)+g(0)=1+1=2\neq1\]

OpenStudy (turingtest):

the question is not whether the result of our addition is in c, but whether it is in our subspace W

OpenStudy (anonymous):

ohhh ok just digesting that

OpenStudy (turingtest):

if the space is not closed by addition and scalar multiplication it is not a subspace, not matter if the result is in the original space V

OpenStudy (anonymous):

like if it was f(0)=0 then it wld be a subspace right?

OpenStudy (turingtest):

Right

OpenStudy (anonymous):

kkk got it Thanks turing, U r a lifesavor

OpenStudy (turingtest):

that, by the way, is called the null space it is a subspace of every vector space

OpenStudy (anonymous):

oh cool

OpenStudy (turingtest):

sorry it's called 'zero space'*

OpenStudy (turingtest):

null space is something else

OpenStudy (turingtest):

every vector space has at least two subspaces 1)the zero space 2)itself and it must be closed in /itself/ hope that helps

OpenStudy (anonymous):

yes i am sord of getting the hang of it

OpenStudy (turingtest):

glad, good luck :D

OpenStudy (anonymous):

thanks